OpenStudy (anonymous):

I need help on a math project... First, I need to figure out 2x^2+5x-12=0 by factoring. I thought I had finally got this down, but I can't figure it out to save my life. I'm supposed to figure out what will multiple to -12 and add to 5, right?

OpenStudy (anonymous):

first imagine that A=2x^2+5x-12 2A=4x^2+10x-24 2A=4x^2+2*5x-24

OpenStudy (anonymous):

Wait, what? I mean... I think I get what you're saying, but I've never done it like that before... Would I take 12 and 2, and then divide them by two to fit the first equation...?

OpenStudy (anonymous):

Wait, no... That doesn't make sense, sorry

OpenStudy (unklerhaukus):

You could use the quadratic formula: A Quadratic Equation\[ax^2+bx+c=0\]The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

I have to use that for a different part of the assignment. I have to find the answer using a table, the quadratic formula, graphing, and factoring

OpenStudy (anonymous):

2A=4x^2+2*5x-24 2A=(2x+8)(2x-3) A=(x+4)(2x-3) when 2x^2+5x-12=(x+4)(2x-3) you can solve this equation

OpenStudy (anonymous):

well if (ax-b)(cx-d)=0 then we can say that ax-b=0 then ax=b then x=b/a .the second answer is cx-d=0 then cx=d then x=d/c for example:(x-4)=0 so x=4

OpenStudy (anonymous):

Okay, I've been working on it... Would the final answer(s) be x = 4 and x = -1.5?

OpenStudy (anonymous):

Wait, no. Reverse those. -4 and 1.5. Or am I wrong..?

OpenStudy (anonymous):

YES.the answers are -4 and 1.5.you can test these answers by plugging them in your equation.

OpenStudy (anonymous):

Yes! Ugh, thank you. I can finally finish!

OpenStudy (anonymous):

\[ 2x^2+5x-12=0\]\[x^2+5x-24=0\]\[(x-8)(x+3)=0\]\[\left(x-\frac{8}{2}\right)\left(x-\frac{3}{2}\right)=0\]Reduce what you can and cross multiply the denominator of the other one over to the variable and reduce your factor.\[(x-4)(2x-3)=0\]Now set each of the factors equal to zero and solve.